# Few-shot Image Generation via Information Transfer from the Built Geodesic Surface

### 03 Jan 2024 v1

@@ 3.2 Interpolation and Regularization (I&\&&R)
The importance of observing the latent space interpolation is to mitigate the potential occurrence of blurry and ”stairlike” interpolation outcomes. Smooth latent space interpolation is an important property of generative models that disproves overfitting and allows synthesis of novel data samples kong_few-shot_2022 . The interpolation set in the latent space $\{z^{\prime}_{1},z^{\prime}_{2},...,z^{\prime}_{k}\}$, predefined as $Z_{inp}(z^{\prime}_{1},z^{\prime}_{k})$, which is obtained by linear interpolating two random latent vectors $z^{\prime}_{1},z^{\prime}_{k}\sim p(z)$. Subsequently, we can obtain a set of interpolated images through $\mathcal{G}(Z_{inp})$ with FAGS employed during the training of $\mathcal{G}$. However, the changes in the appearances of $\mathcal{G}(Z_{inp})$ are discontinuous, causing the occurrence of blurriness in the intermediate interpolations. For instance, Figure 3 visually portrays one of the $\mathcal{G}(Z_{inp})$ through the trained StyleGAN2 karras_analyzing_2020 with FAGS.
To alleviate the blurry appearance and generate more realistic images, we present an interpolation strategy. To supervise $\mathcal{G}(Z_{inp})$, the original adversarial loss goodfellow_generative_2020 can be rewriten based on Formula 1 and 2 for both generator $\mathcal{G}$ and discriminator $\mathcal{D}$, as follows:
@@
To further avoid “stairlike” phenomenon, we propose a distance regularization strategy. We adopt the KL-Divergence loss kullback1951information to enforce equal pairwise distances between features of interpolated images, namely $L_{dr}$. Average pooling is used to aggregate the feature map for minimizing the computational time and space, reducing both the $H$ and $W$ down to 1/4 of the original size. The proposed strategy penalizes the generator $\mathcal{G}$ from merely memorizing real samples and helps to generate smooth images during latent space interpolation.
The two strategies consist of a module, named as Interpolation and Regularization (I&R), and shown in Figure4. Algorithm 1 outlines the pseudo-code of the I&R module.
Algorithm 1 Pseudo-code of I&R module in a PyTorch-like style.
@@
6:$\rm pred=Discriminator(inp\_imgs)$
7:$\rm L\_inp=log(pred\_imgs).mean()$
8:
9:#$\rm\#\ \ Regularization$
10:$\rm dist()=L2\_distance()$
11:$\rm inp\_feats=AdaptiveAvgPool2d(inp\_feats)\ \ \ \#\ k\times c\times h\times w\rightarrow k\times c\times h/4\times w/4$
[…]
13:$\rm q\_dist=cat([ones(k-1),Tensor([k-1])])\ \ \ \#\ k\times 1$
14:$\rm L\_dr=KLDivLoss(log\_softmax(feats\_dist),normalize(q\_dist))$
3.3 Final Optimization Function
The final optimization function $L^{G}$ for generator $\mathcal{G}$ and $L^{D}$ for discriminator $\mathcal{D}$ are defined as follows:
@@ 3.3 Final Optimization Function
4 Experiments and Discussion
4.1 Implementation details
We employ StyleGAN2 karras_analyzing_2020 with MixDL kong_few-shot_2022 as our backbone architecture without using ADA karras2020training to deal with the extreme few-shot scenario. We set the parameters of Formula 14 and 15 as follows: $\lambda_{1}=0.8,\lambda_{2}=1.25,$ and $\lambda_{3}=0.8$. Both batch size and interpolation size are configured as 4 to ensure compatibility with a single Nvidia GeForce RTX 3090 (24GB) for training.
Our experiments involve qualitative and quantitative comparisons with several models, including N-div liu2019normalized , MSGAN mao2019mode , DistanceGAN benaim2017one , StyleGAN2 (SG2) karras_analyzing_2020 , StyleGAN2+ADA (SG2A) karras2020training , FastGAN (FG) liu_towards_2021 , and MixDL (MDL) kong_few-shot_2022 .
We experiment on multiple datasets, including Amedeo Modigliani paintings yaniv2019face , Landscape drawings ojha_few-shot_2021 , Animal-Face Dog si2011learning , Face sketches wang2008face , Anime face liu_towards_2021 , Pokemon liu_towards_2021 , FFHQ karras_style-based_2019 , and CelebA liu2018large . Specifically, Amedeo Modigliani paintings and Landscape drawings are 10-shot datasets. For the remaining datasets, which contain more than 10 images, 10 images are randomly chosen from each to construct the 10-shot subsets. All images in our experiments are of resolution 256 $\times$ 256.
4.2 Qualitative Comparison
@@
In Figure 5, the results of various methods are presented on the Amedeo Modigliani paintings and Landscape drawings datasets. It is important to note that all the models are trained from scratch without auxiliary information produced, except for FastGAN liu_towards_2021 , which utilized a pre-trained VGG for calculating perceptual loss zhang_unreasonable_2018 . From Figure 5, the generated samples of StyleGAN2 karras_analyzing_2020 show blurriness and overfitting on both 10-shot datasets. FastGAN liu_towards_2021 produces comparable results to our method on the Amedeo Modigliani paintings dataset, but demonstrates inferior quality on the Landscape drawings dataset. FastGAN merely weighted add two training samples, as depicted in the generated sample circled in red in Figure 5. MixDL-generated images kong_few-shot_2022 excel in terms of diversity but lack in fidelity compared to other methods. For example, the face shape is distorted in the generated sample circled in blue in Figure 5. Notably, ITBGS stands out by generating images with both sufficient fidelity and diversity on both 10-shot datasets. The achievement is attributed to the natural fusion of visual elements, such as shapes, colors, textures, from two or more real images.
@@
In Figure 6, 7, 8 and 9, we provide a comparison of the interpolated samples generated with some methods across various 10-shot datasets. Notably, FastGAN liu_towards_2021 achieves comparable generative results to our method on the Face sketches dataset. However, FastGAN-generated interpolated samples exhibit the “stairlike” phenomenon from column f to column g in Figure 6 (Amedeo Modigliani). Also, defects occur in column c of Figure 9. MixDL kong_few-shot_2022 excels in producing smooth semantic interpolations but comes at the cost of reduced fidelity. For instance, the mouth and the mountain peaks appear strange in the red-circled sections of Figure 6 and 7, respectively. StyleGAN2+ADA karras2020training also demonstrates a similar trade-off of fidelity and diversity on Amedeo Modigliani and Anime Face dataset. In contrast, the original StyleGAN2 karras_analyzing_2020 generates interpolated images with acceptable fidelity but exhibits the “stairlike” phenomenon, as depicted from column e to column f in Figure 6 and 8. ITBGS shows the capability of achieving smooth latent space interpolation across all 10-shot datasets while maintaining sufficient fidelity. Smooth latent space interpolation underlines the effectiveness of ITBGS in balancing fidelity and diversity for image generation.
@@
The quality of results in generating real-world images serves as a valuable metric for evaluating the effectiveness of generative models. The experiments on real-world face datasets FFHQ karras_style-based_2019 , as depicted in Figure 10, demonstrates a relatively satisfactory ability to combine facial features from two or more faces. The ability of combination is particularly evident in the seamless blending of hairstyles, beards, and other facial attributes. Notably, FastGAN liu_towards_2021 also produces decent results on the FFHQ dataset, while the other comparative methods struggle to generate real-world facial images.
4.3 Quantitative comparison
Table 1: Quantitative results on 10-shot image generation task. The best and the second best scores are in bold and underlined, respectively.
@@ 4.3 Quantitative comparison
N-Divliu2019normalized 175.4 0.425 150.4 0.632 / / / / / / / /
MSGANmao2019mode 138.6 0.536 165.7 0.630 / / / / / / / /
[…]
StyleGAN2+ADAkarras2020training 282.3 0.473 342.0 0.539 341.3 0.469 216.3 0.538 207.7 0.498 278.5 0.413
[…]
In the evaluation, we employ the Fréchet Inception Distance (FID) heusel2017gans as well as the pairwise Learned Perceptual Image Patch Similarity (LPIPS) zhang_unreasonable_2018 as metrics. FID is computed against the few-shot dataset, while LPIPS is calculated between generated samples. Lower FID values and higher LPIPS values are indicative of higher image quality and diversity, respectively.
Table 2: Quantitative results on FFHQ and CelebA. The best and the second best scores are in bold and underlined, respectively.
@@
[…]
Tables 1 and 2 present the quantitative results of the compared methods on various 10-shot datasets. We use consistent parameter settings across different datasets and conduct no domain-specific fine-tuning. The tables show that our method achieves the best or near-best results in terms of both FID and LPIPS. Assessing the capabilities of generative model requires considering both fidelity and diversity, which are reflected by these metrics. While the scores of our method are slightly lower than FastGAN liu_towards_2021 on Pokemon and Landscape datasets, it still ranks the second-best position. Moreover, as demonstrated in Figure 10, the generated samples by our method and FastGAN on a real-world face dataset yield comparable visual quality. However, when examining the quantitative metrics in Table. 2, our method outperforms FastGAN. While FastGAN has a slightly lower FID, our method excels in terms of diversity.
4.4 Ablation Study
4.4.1 Effect of the FAGS module
@@
Table 3: Quantitative ablation on FAGS module.
@@
To validate the effectiveness of the proposed FAGS module, we conducted ablation experiments. As depicted in Figure 11, the FAGS with $L_{g}$ significantly enhances the visual quality of the generated samples. The method of RSSA xiao_few_2022 with self-correlation consistency loss $L_{scc}$, which removes the pre-trained generative model, use the training images directly as the source domain and transfer their information to the target generator. However, due to the limited information in the few-shot source domain, the method leads to blurry images. We address these problems by introducing the FAGS module to create a pseudo-source domain by building the Geodesic surface. However, if we replace the loss function from $L_{scc}$ to others, like the $-\ell$1 loss, it results in inadequate fidelity. Similar conclusions can be drawn from the quantitative results presented in Table 3.
4.4.2 Effect of the I&\&&R module

### 03 Mar 2024 v2

@@ III-B Interpolation and Regularization (I&\&&R)
The importance of observing the latent space interpolation is to mitigate the potential occurrence of blurry and ”stairlike” interpolation outcomes. Smooth latent space interpolation is an important property of generative models that disproves overfitting and allows synthesis of novel data samples [10]. The interpolation set in the latent space $\{z^{\prime}_{1},z^{\prime}_{2},...,z^{\prime}_{k}\}$, predefined as $Z_{inp}(z^{\prime}_{1},z^{\prime}_{k})$, which is obtained by linear interpolating two random latent vectors $z^{\prime}_{1},z^{\prime}_{k}\sim p(z)$. Subsequently, we can obtain a set of interpolated images through $\mathcal{G}(Z_{inp})$ with FAGS employed during the training of $\mathcal{G}$. However, the changes in the appearances of $\mathcal{G}(Z_{inp})$ are discontinuous, causing the occurrence of blurriness in the intermediate interpolations. For instance, Figure 3 visually portrays one of the $\mathcal{G}(Z_{inp})$ through the trained StyleGAN2 [44] with FAGS.
To alleviate the blurry appearance and generate more realistic images, we present an interpolation strategy. To supervise $\mathcal{G}(Z_{inp})$, the original adversarial loss [2] can be rewriten based on Formula 1 and 2 for both generator $\mathcal{G}$ and discriminator $\mathcal{D}$, as follows:
@@
To further avoid “stairlike” phenomenon, we propose a distance regularization strategy. We adopt the KL-Divergence loss [45] to enforce equal pairwise distances between features of interpolated images, namely $L_{dr}$. Average pooling is used to aggregate the feature map for minimizing the computational time and space, reducing both the $H$ and $W$ down to 1/4 of the original size. The proposed strategy penalizes the generator $\mathcal{G}$ from merely memorizing real samples and helps to generate smooth images during latent space interpolation.
The two strategies consist of a module, named as Interpolation and Regularization (I&R), as shown in Figure 4. Algorithm 1 outlines the pseudo-code of the I&R module in a PyTorch-like style.
Algorithm 1 Pseudo-code of I&R module.
@@
5:$\rm inp\_imgs,inp\_feats=Generator(Z\_inp)$
6:#$\#\ inp\_feats:k\times c\times h\times w$
7:$\rm pred=Discriminator(inp\_imgs)$
8:$\rm L\_inp=log(pred\_imgs).mean()$
9:
10:#$\rm\#\ \ Regularization$
11:$\rm dist()=L2\_distance()$
12:inp_feats=AdaptiveAvgPool2d($\rm inp\_feats=AdaptiveAvgPool2d(inp\_feats)$
13:#$\#\ k\times c\times h\times w\rightarrow k\times c\times h/4\times w/4$
@@
15:$\rm feats\_dist=dist(inp\_feats,inp\_feats\_temp)$
16:$\rm q\_dist=cat([ones(k-1),Tensor([k-1])])$
17:$\rm L\_dr=KLDivLoss(feats\_dist,q\_dist)$
III-C Final Optimization Function
The final optimization function $L^{\mathcal{G}}$ for generator $\mathcal{G}$ and $L^{\mathcal{D}}$ for discriminator $\mathcal{D}$ are defined as follows:
@@ III-C Final Optimization Function
IV Experiments and Discussion
IV-A Implementation details
We employ StyleGAN2 [44] with MixDL [10] as our backbone architecture without using ADA [8] to deal with the extreme few-shot scenario. We set the parameters of Formula 14 and 15 as follows: $\lambda_{1}=0.8,\lambda_{2}=1.25,$ and $\lambda_{3}=0.8$. Both batch size and interpolation size are configured as 4 to ensure compatibility with a single Nvidia GeForce RTX 3090 (24GB) for training.
Our experiments involve qualitative and quantitative comparisons with several models, including N-div [46], MSGAN (MG) [47], DistanceGAN (DG) [48], StyleGAN2 (SG2) [44], StyleGAN2+ADA (SG2A) [8], FastGAN (FG) [9], and MixDL (MDL) [10].
We experiment on multiple datasets, including Amedeo Modigliani paintings [49], Landscape drawings [6], Animal-Face Dog [50], Face sketches [51], Anime face [9], Pokemon [9], FFHQ [52], and CelebA [53]. Specifically, Amedeo Modigliani paintings and Landscape drawings are 10-shot datasets. For the remaining datasets, which contain more than 10 images, 10 images are randomly chosen from each to construct the 10-shot subsets. All images in our experiments are of resolution 256 $\times$ 256.
IV-B Qualitative Comparison
@@
In Figure 5, the results of various methods are presented on the Amedeo Modigliani paintings and Landscape drawings datasets. It is important to note that all the models are trained from scratch without auxiliary information produced, except for FastGAN [9], which utilized a pre-trained VGG for calculating perceptual loss [54]. From Figure 5, the generated samples of StyleGAN2 [44] show blurriness and overfitting on both 10-shot datasets. FastGAN [9] produces comparable results to our method on the Amedeo Modigliani paintings dataset, but demonstrates inferior quality on the Landscape drawings dataset. FastGAN merely weighted add two training samples, as depicted in the generated sample circled in red in Figure 5. MixDL-generated images [10] excel in terms of diversity but lack in fidelity compared to other methods. For example, the face shape is distorted in the generated sample circled in blue in Figure 5. Notably, ITBGS stands out by generating images with both sufficient fidelity and diversity on both 10-shot datasets. The achievement is attributed to the natural fusion of visual elements, such as shapes, colors, textures, from two or more real images. We hypothesize that the features generated on the Geodesic surface, constructed in the Pre-Shape Space using extracted image features from the dataset, represent a natural integration of from multiple image features. Consequently, these generated features facilitate the generation of images that exhibit a more natural fusion of visual elements.
@@
In Figure 6, 7, 8 and 9, we provide a comparison of the interpolated samples generated with some methods across various 10-shot datasets. Notably, FastGAN [9] achieves comparable generative results to our method on the Face sketches dataset. However, FastGAN-generated interpolated samples exhibit the “stairlike” phenomenon from column f to column g in Figure 6 (Amedeo Modigliani). Also, defects occur in column c of Figure 9. MixDL [10] excels in producing smooth semantic interpolations but comes at the cost of reduced fidelity. For instance, the mouth and the mountain peaks appear strange in the red-circled sections of Figure 6 and 7, respectively. StyleGAN2+ADA [8] also demonstrates a similar trade-off of fidelity and diversity on Amedeo Modigliani and Anime Face dataset. In contrast, the original StyleGAN2 [44] generates interpolated images with acceptable fidelity but exhibits the “stairlike” phenomenon, as depicted from column e to column f in Figure 6 and 8. ITBGS shows the capability of achieving smooth latent space interpolation across all 10-shot datasets while maintaining sufficient fidelity. Smooth latent space interpolation underlines the effectiveness of ITBGS in balancing fidelity and diversity for image generation.
@@
The quality of results in generating real-world images serves as a valuable metric for evaluating the effectiveness of generative models. The experiments on real-world face datasets FFHQ [52], as depicted in Figure 10, demonstrate a relatively satisfactory ability to combine facial features from two or more faces. The ability of combination is particularly evident in the seamless blending of hairstyles, beards, and other facial attributes. Notably, FastGAN [9] also produces decent results on the FFHQ dataset, while the other comparative methods struggle to generate real-world facial images.
IV-C Quantitative comparison
TABLE I: Quantitative results on 10-shot image generation task. The best and the second best scores are in bold and underlined, respectively.
@@ IV-C Quantitative comparison
N-Div[46] 175.4 0.425 150.4 0.632 / / / / / / / /
MSGAN[47] 138.6 0.536 165.7 0.630 / / / / / / / /
DistanceGAN[48] 84.1 0.543 102.6 0.678 / / / / / / / /
StyleGAN2[44] 213.9 0.407 312.9 0.549 188.4 0.476 68.6 0.649 210.3 0.531 261.9 0.475
StyleGAN2+ADA[8] 282.3 0.473 342.0 0.539 341.3 0.469 216.3 0.538 207.7 0.498 278.5 0.413
MixDL[10] 140.9 0.529 291.1 0.701 137.9 0.396 205.2 0.643 183.3 0.698 231.2 0.499
FastGAN[9] 150.5 0.393 65.1 0.671 112.4 0.437 108.3 0.615 83.8 0.689 203.3 0.554
FastGAN+FAGS 123.2 0.304 54.5 0.679 97.8 0.292 98.9 0.588 82.6 0.699 200.8 0.420
Ours 72.5 0.538 95.0 0.713 57.7 0.485 113.9 0.647 90.7 0.677 208.1 0.552
In the evaluation, we employ the Fréchet Inception Distance (FID) [55] as well as the pairwise Learned Perceptual Image Patch Similarity (LPIPS) [54] as metrics. FID is computed against the few-shot dataset, while LPIPS is calculated between generated samples. Lower FID values and higher LPIPS values are indicative of higher image quality and diversity, respectively.
TABLE II: Quantitative results on FFHQ and CelebA. The best and the second best scores are in bold and underlined, respectively.
@@
FastGAN+FAGS 220.9 0.448 67.3 0.554
Ours 130.9 0.617 91.3 0.570
Tables I and II present the quantitative results of the compared methods on various 10-shot datasets. We use consistent parameter settings across different datasets and conduct no domain-specific fine-tuning. The tables show that our method achieves the best or near-best results in terms of both FID and LPIPS. Assessing the capabilities of generative model requires considering both fidelity and diversity, which are reflected by these metrics. While the performance metrics of our method are slightly lower than FastGAN [9] on Pokemon and Landscape datasets, it still ranks the third-best position. On the real-world face dataset, the generated samples by our method and FastGAN yield comparable visual quality, as demonstrated in Figure 10. However, when examining the quantitative metrics in Table. II, our method outperforms FastGAN. While FastGAN has a slightly lower FID, our method excels in terms of diversity. Furthermore, the integration of the FAGS module into FastGAN results in enhanced performance on some of the datasets, including Pokemon, Landscapes, and CelebA datasets.
IV-D Ablation Study
TABLE III: Quantitative ablation on the proposed modules.
Backbone FAGS I R Landscapes
FID(↓) LPIPS(↑)
StyleGAN2 × × × 210.3 0.531
× × 206.9(-3.4) 0.629(+0.098)
× 182.3(-24.6) 0.669(+0.04)
90.7(-91.6) 0.677(+0.008)
FastGAN × × × 83.8 0.689
× × 82.6(-1.2) 0.699(+0.01)
97.5(+14.9) 0.679(-0.02)
We conduct quantitative ablation experiments with the proposed modules. As shown in Table III, the integration of the FAGS module enhances the performance metrics for both backbones: StyleGAN2 and FastGAN. The deployment of the I&R module contributes to a further performance increment for StyleGAN2, since adding only FAGS may results in blurry generated images, as shown in Figure 3. However, for FastGAN, adding the I&R module results in a inferior performance compared to the sole addition of the FAGS module, which suggests that the I&R module may be backbone-sensitive.
@@ IV-D Ablation Study
TABLE IV: Quantitative ablation on FAGS module.
@@
IV-D1 Effect of the FAGS module
To validate the effectiveness of the proposed FAGS module, we conducted ablation experiments. As depicted in Figure 11, the FAGS with $L_{g}$ significantly enhances the visual quality of the generated samples. The method of RSSA [7] with self-correlation consistency loss $L_{scc}$, which removes the pre-trained generative model, use the training images directly as the source domain and transfer their information to the target generator. However, due to the limited information in the few-shot source domain, the method leads to blurry images. We address these problems by introducing the FAGS module to create a pseudo-source domain by building the Geodesic surface. However, if we replace the loss function from $L_{scc}$ to others, like the $-\ell$1 loss, it results in inadequate fidelity. Similar conclusions can be drawn from the quantitative results presented in Table IV.
IV-D2 Effect of the I&\&&R module
@@ V Conclusion
Acknowledgments
This research is sponsored by National Natural Science Foundation of China (Grant No. 52273228), Key Research Project of Zhejiang Laboratory (No. 2021PE0AC02), Key Program of Science and Technology of Yunnan Province (202302AB080022), the Project of Key Laboratory of Silicate Cultural Relics Conservation (Shanghai University), Ministry of Education (No. SCRC2023ZZ07TS).